Using the information in the table below, calculate the cost breakdown per item and the monthly...

  1. Using the information in the table below, calculate the cost breakdown per item and the monthly cost for the release of information (ROI) services at Community Hospital. Round wages to whole number.

Item

Cost per Month

Postage

$675

Service contract (includes copier)

$250

Equipment (includes copies)

$125

Supplies (includes toner, printer cartridges, paper)

$95

Wages

$12.00 per hour

Monthly cost
Answer #1

Average requests per month is 540

a. Postage cost per month = $ 675

So the Postage cost request per one month= 675/540 = $ 1.25

b. Service contract (includes copier) cost per month = $ 250

So the Service contract cost request per one month= 250/540 = $ 0.46

c. Equipment (includes copies) cost per month = $ 125

So the Equipment cost request per one month= 125/540 = $ 0.2

d. Supplies (includes toner, printer cartridges, paper) cost per month = $ 95

So the supplies cost request per one month= 95/540 = $ 0.17

e. Wages 12 $ per hour

12 X 2080 hours= 24960 $ annual salary

So the monthly salary is = 24960/12 = 2080

So for one request per month= 2080/540 = $ 3.85

Monthly cost = $ 5. 93 (1.25 + 0.46 + 0.20 + 0.17 + 3. 85)